OpenStudy (anonymous):

The two ends of a train moving with uniform acceleration pass a certain point with velocities u and v.The velocity with which the middle point of the train passes the same point is?

OpenStudy (nikvist):

\[v^2=u^2+2as\rightarrow s=\frac{v^2-u^2}{2a}\] \[w^2=u^2+as\rightarrow w=\sqrt{u^2+\frac{v^2-u^2}{2}}\]